OpenStudy (anonymous):

MEADL FOR RIGHT ANSWER The temperature was 65 degrees at daybreak. Then it dropped two degrees per hour until dusk. If x is the number of hours since daybreak, this equation represents the temperature during the day: y = -2x + 65 OR y = 65 - 2x. Use the equation to predict the temperature six hours later.

OpenStudy (anonymous):

You can't do anything. First of all the equations don't make sense. There is nothing to combine.

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Nvr mind

OpenStudy (anonymous):

a. 53 b. 63 c. 61 d. 59

OpenStudy (anonymous):

Okay what is the formula for calculating temperature?

OpenStudy (anonymous):

Ok, so let's use y = 65 - 2x

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

Since x is the number of hours, and it asks what the temperature will be in 6 hours, you replace x with 6. So the equation is y = 65 - 2 x 6

OpenStudy (anonymous):

Ok, so what is 2 x 6?

OpenStudy (anonymous):

How do u know that x = 6?

OpenStudy (anonymous):

12, 65-12 = 53

OpenStudy (anonymous):

i missed that part ;lol

OpenStudy (anonymous):

Yep, that's your answer, @hhmh4 !

OpenStudy (anonymous):

thanxs guys

OpenStudy (anonymous):

I see how u got 6 now

OpenStudy (anonymous):

Yeah.

OpenStudy (anonymous):

Np.